# Observational Probes of the Neutron Star Equation of State with Hyperons, Bosonic Dark Matter, and Quark Matter

### v2

@@ 5 Bayesian Inference from Neutron Star Observations
To constrain the value of the S mass in our BA, we incorporate a set of astrophysical measurements leading to M-R and tidal deformability constraints presented in the previous section. Additionally, we examine the influence of individual subsets of these constraints on our EOS model estimates to assess the impact of the S mass.
@@ 5.1 Basic Observational Dataset
    Precise mass measurement of PSR J0348+0432 in a white dwarf binary (Antoniadis et al., 2013):
    $M=2.01^{+0.04}_{-0.04}\leavevmode\nobreak\ M_{\odot}.$
@@
    NICER measurements for mass and radius for PSR J0740+6620 from Dittmann et al. (2024b), which is recent refinement of Miller et al. (2021):
    $M=2.08^{+0.07}_{-0.07}\leavevmode\nobreak\ M_{\odot},\quad R=12.92^{+2.09}_{-1.13}\leavevmode\nobreak\ \mathrm{km}.$
@@
    NICER measurements for PSR J0030+0451 from Vinciguerra et al. (2024), which is recent refinement of Riley et al. (2019):
    $M=1.44^{+0.15}_{-0.14}\leavevmode\nobreak\ M_{\odot},\quad R=13.02^{+1.24}_{-1.06}\leavevmode\nobreak\ \mathrm{km}.$
@@
    Mass and radius of PSR J0437–4715 (Choudhury et al., 2024b):
    $M=1.418^{+0.037}_{-0.037}\leavevmode\nobreak\ M_{\odot},\quad R=11.36^{+0.95}_{-0.63}\leavevmode\nobreak\ \mathrm{km}.$
@@
    Additionally, we consider the tidal deformability constraint from the GW170817 event (Abbott et al., 2018):
@@ 5.2 Extended Dataset
Additionally we test the results by including two extreme measurements:
@@
    High-mass BW pulsar PSR J0952-0607 (Romani et al., 2022):
    $M=2.35^{+0.17}_{-0.17}\leavevmode\nobreak\ M_{\odot}.$
@@
    Low-mass ultra compact object HESS J1731-347 (Doroshenko et al., 2022):
    $M=0.77^{+0.20}_{-0.17}\leavevmode\nobreak\ M_{\odot},\quad R=10.04^{+0.86}_{-0.78}\leavevmode\nobreak\ \mathrm{km}.$
Furthermore, we examine EOSs in light of the newly reported mass and radius measurements for the rotation-powered millisecond pulsar PSR J0614-3329, based on NICER observations, which indicate that it is a moderately massive and relatively compact object, possessing the same radius as HESS J1731-347 but nearly twice the mass:
@@
    Moderate-mass compact millisecond pulsar PSR J0614-3329 (Mauviard et al., 2025b):
    $M=1.44^{+0.06}_{-0.07}\leavevmode\nobreak\ M_{\odot},\quad R=10.29^{+1.01}_{-0.86}\leavevmode\nobreak\ \mathrm{km}.$
@@ 5.3 Bayesian Analysis Framework
$p(D_{M}|m_{\text{S}}=F_{\mathcal{N}}(M_{\max}(m_{\text{S}});\mu_{M},\sigma_{M}),$ (17)
@@
For M-R constraints, we evaluate the maximum likelihood approach:
@@
The density functions $f_{MR}$ are obtained via kernel density estimation (KDE) using public datasets from Zenodo: (Vinciguerra et al., 2023; Dittmann et al., 2024a; Choudhury et al., 2024a; Doroshenko et al., 2023; Mauviard et al., 2025a).
Similarly, the GW170817 constraint is evaluated as
@@
Finally, the Bayesian evidence is computed as:
@@
Note that we prefer the maximum likelihood approach over numerical marginalization, as employed, for example, in Alvarez-Castillo et al. (2016); Shahrbaf et al. (2023); Ayriyan et al. (2025), because marginalizing over central density can introduce noticeable numerical inaccuracies when the EOS stiffness varies significantly within the target region. A fixed step in central density may lead to significant uneven sampling in observable space (e.g., mass and radius) across different EOSs, potentially biasing the results. We observed this effect in the region of the most probable NS masses, where our EOS exhibits substantial variation in stiffness. In this regime, numerical marginalization using M-R constraints from Vinciguerra et al. (2023), Choudhury et al. (2024b) and Mauviard et al. (2025b) becomes less reliable for our case. In contrast, the maximum likelihood method directly identifies the best-fitting models without relying on sensitive numerical integration.
@@ 5.4 Bayesian Analysis Results
Incorporating the exotic compact objects HESS J1731–347 and PSR J0952–0607 in Fig. 8 further reinforces this trend. In particular, the lightness and high compactness of HESS J1731–347 strongly favor S masses $m_{\text{S}}\lesssim 1900$ MeV 8(b). While our model intersects only the higher right edge of the 95% confidence region inferred for HESS J1731–347 by Doroshenko et al. (2022). This happens because the constraint yields a higher weighted likelihood difference for different values of $m_{S}$.
It is worth noting that, due to the unexpectedly high compactness of this source, we expect that future refinements may lead to significant changes in its estimated mass and radius in favor of less compactness, despite this, various hypotheses are now being put forward about the nature of this ultra-compact object. Several studies have proposed the nature of such light and ultra-compact objects, it could be strange stars (Di Clemente et al., 2024; Horvath et al., 2023; Yuan & Zhou, 2025), hybrid stars with a deconfined quark core (Ayriyan et al., 2025), or even third family compact stars (Ayriyan et al., 2025; Alvarez-Castillo, 2025). On the basis of the obtained results, we alternatively suggest that the presence of DM in such objects is a plausible hypothesis.
The final figure 8(c) illustrates the results of the analysis including all the constraints on mass and radius (I)-(VII). The dominant contribution to the posterior distribution arises from constraints in the region of canonical masses, as well as from the ultra-compact object, resulting in an even stronger preference for lower masses of S.
@@
Figure 8: Posterior for additional set of constraints: (VI) high-mass BW pulsar PSR J0952-0607 (Romani et al., 2022) and (VII) ultra-compact HESS J1731-347 (Doroshenko et al., 2022).
@@
Figure 9: Posterior distribution with new NICER constraints (VIII) from the millisecond pulsar PSR J0614–3329 (Mauviard et al., 2025b)
@@ 6 Summary and conclusions
Motivated by the theoretical existence of a stable sexaquark configuration S=uuddss, with a mass below $2054$ MeV, we have explored its role as a bosonic dark matter (DM) candidate within the core of neutron stars (NSs). Using the DD2Y-T model to describe the hadronic phase, incorporating both hyperons and the sexaquark (S) as a bosonic DM component, we modeled the superconducting quark matter (QM) phase with the non-local NJL (nlNJL) model. A smooth crossover transition between these two phases was constructed using the Replacement Interpolation Construction method. We show that all hybrid stars, regardless of the mass of the S, contain approximately $16\%$ DM.
Our results show that the inclusion of S particles with a weak repulsive interaction, modeled via a positive mass shift, leads to a softening of the equation of state (EOS). This softening plays a critical role in addressing the tension between stiff hadronic EOSs and tidal deformability constraints, particularly around the canonical $1.4M_{\odot}$ mass range. In contrast to purely hadronic or hybrid models without DM, the resulting hybrid stars, comprising a deconfined QM core and a hypernuclear outer layer enriched with S DM, satisfy all current observational constraints on NS properties.
Incorporating recent astrophysical observations, we find that for $m_{\text{S}}\leq 1930$ MeV, the mass–radius (M-R) relations of the resulting hybrid stars not only remain consistent with existing measurements but also lie within the $68\%$ credible region from NICER for PSR J0437-4715 around $1.4M_{\odot}$. Moreover, the hybrid star configurations with the two lowest S masses used in this work agree with the constraints inferred from the low-mass ultra-compact object HESS J1731–347. For $20401885<m_{\text{S}}<$ 2040 MeV, the tidal deformability $\Lambda_{1.4}$ remains below the observational upper limit of 580, further supporting the viability of our EOS and indicating the highest allowed mass for S particles. It is worth noting that without the presence of DM, the considered hadronic (DD2Y-T) and hybrid (RIC_DD2Y-T) models are not consistent with $\Lambda_{1.4}$. In addition, by employing the recently updated NICER M-R results for PSR J0030+0451, we showed that the EOSs including S DM are compatible with the most probable observational values, i.e., the $68\%$ credible region, while the hybrid model without S and the pure hadronic EOS are only consistent with the $95\%$ region.
Furthermore, to place the most stringent constraint on the S mass while accounting for all available observational data, we have incorporated the latest NICER M-R measurements of PSR J0614–3329, which became available prior to the submission of this study. We find that, unlike the EOSs without the S component, all models that include DM favor the $95\%$ credible region, with lighter S particles being particularly preferred. This noteworthy result provides strong support for our model, in which S particles are produced inside NSs, coexisting with other possible internal structures, including hyperons and QM.
@@
To investigate the parameter space of S and constrain its mass more precisely, we performed a Bayesian analysis that incorporates all key observational constraints within a consistent statistical framework. This includes all available M-R measurements by the NICER telescope and the tidal deformability constraint from the GW170817 event by LIGO/Virgo detectors. Additionally, we consider two astrophysical extremes: HESS J1731–347, representing the most compact (lightest) known NS, and PSR J0952–0607, the most massive NS observed to date. These sources provide critical boundary conditions for constraining the EOS across the full density range.
@@
While our study does not claim that the presence of S particles is required to explain NS properties, it provides compelling evidence that their inclusion leads to improved consistency with current observational data. Future observations, particularly those sensitive to the thermal and cooling behavior of NSs, may offer further insight into the role of such exotic particles in the interiors of compact stars.

### v3

@@ 5 Bayesian inference from neutron star observations
To constrain the value of the S mass in our BA, we incorporated a set of astrophysical measurements leading to M-R and tidal deformability constraints presented in the previous section. Additionally, we examined the influence of individual subsets of these constraints on our EOS model estimates to assess the impact of the S mass.
@@ 5.1 Basic observational dataset
    Precise mass measurement of PSR J0348+0432 in a white dwarf binary (Antoniadis and others, 2013):
    $M=2.01^{+0.04}_{-0.04}\penalty 10000\ M_{\odot}.$
@@
    The NICER measurements for mass and radius for PSR J0740+6620 from Dittmann and others (2024), which is a recent refinement of Miller and others (2021):
    $M=2.08^{+0.07}_{-0.07}\penalty 10000\ M_{\odot},\quad R=12.92^{+2.09}_{-1.13}\penalty 10000\ \mathrm{km}.$
@@
    The NICER measurements for PSR J0030+0451 from Vinciguerra and others (2024), which is a recent refinement of Riley and others (2019):
    $M=1.44^{+0.15}_{-0.14}\penalty 10000\ M_{\odot},\quad R=13.02^{+1.24}_{-1.06}\penalty 10000\ \mathrm{km}.$
@@
    The mass and radius of PSR J0437-4715 (Choudhury and others, 2024):
    $M=1.418^{+0.037}_{-0.037}\penalty 10000\ M_{\odot},\quad R=11.36^{+0.95}_{-0.63}\penalty 10000\ \mathrm{km}.$
@@
    Additionally, we considered the tidal deformability constraint from the GW170817 event (Abbott and others, 2018):
@@ 5.2 Extended dataset
Additionally, we tested the results by including two extreme measurements:
@@
    The high-mass BW pulsar PSR J0952-0607 (Romani et al., 2022):
    $M=2.35^{+0.17}_{-0.17}\penalty 10000\ M_{\odot}.$
@@
    The low-mass ultra compact object HESS J1731-347 (Doroshenko et al., 2022):
    $M=0.77^{+0.20}_{-0.17}\penalty 10000\ M_{\odot},\quad R=10.04^{+0.86}_{-0.78}\penalty 10000\ \mathrm{km}.$
Furthermore, we examined EOSs in light of the newly reported mass and radius measurements for the rotation-powered millisecond pulsar PSR J0614-3329, based on NICER observations, which indicate that it is a moderately massive and relatively compact object, possessing the same radius as HESS J1731-347 but nearly twice the mass:
@@
    Moderate-mass compact millisecond pulsar PSR J0614-3329 (Mauviard and others, 2025):
    $M=1.44^{+0.06}_{-0.07}\penalty 10000\ M_{\odot},\quad R=10.29^{+1.01}_{-0.86}\penalty 10000\ \mathrm{km}.$
@@ 5.3 Bayesian analysis framework
$p(D_{M}|m_{\text{S}})=F_{\mathcal{N}}(M_{\max}(m_{\text{S}});\mu_{M},\sigma_{M}),$ (17)
@@
For M-R constraints, we evaluated the maximum likelihood approach:
@@
The density functions $f_{MR}$ are obtained via kernel density estimation (KDE) using public datasets from Zenodo (Vinciguerra et al., 2023; Dittmann et al., 2024; Choudhury et al., 2024; Doroshenko et al., 2023; Mauviard et al., 2025).
Similarly, the GW170817 constraint was evaluated as
@@
Finally, the Bayesian evidence was computed as
@@
Note that we prefer the maximum likelihood approach over numerical marginalization, as employed, for example, in Alvarez-Castillo et al. (2016); Shahrbaf et al. (2023); Ayriyan et al. (2025), because marginalizing over central density can introduce noticeable numerical inaccuracies when the EOS stiffness varies significantly within the target region. A fixed step in central density may lead to significant uneven sampling in observable space (e.g., mass and radius) across different EOSs, potentially biasing the results. We observed this effect in the region of the most probable NS masses, where our EOS exhibits substantial variation in stiffness. In this regime, numerical marginalization using M-R constraints from Vinciguerra et al. (2023), Choudhury and others (2024) and Mauviard and others (2025) becomes less reliable for our case. In contrast, the maximum likelihood method directly identifies the best-fitting models without relying on sensitive numerical integration.
@@ 5.4 Bayesian analysis results
Incorporating the exotic compact objects HESS J1731–347 and PSR J0952–0607 in Fig. 9 further reinforces this trend. In particular, the lightness and high compactness of HESS J1731–347 strongly favor S masses $m_{\text{S}}\lesssim 1900$ MeV(Fig. 9(b)), while our model intersects only the higher right edge of the 95% confidence region inferred for this object by Doroshenko et al. (2022). This happens because the constraint yields a higher weighted likelihood difference for different values of $m_{S}$.
It is worth noting that, due to the unexpectedly high compactness of this source, we expect that future refinements may lead to significant changes in its estimated mass and radius in favor of less compactness, despite this, various hypotheses are now being put forward about the nature of this ultracompact object. Several studies have proposed the nature of such light and ultracompact objects, it could be strange stars (Di Clemente et al., 2024; Horvath et al., 2023; Yuan and Zhou, 2025), hybrid stars with a deconfined quark core (Ayriyan et al., 2025), or even third family compact stars (Ayriyan et al., 2025; Alvarez-Castillo, 2025). On the basis of the obtained results, we alternatively suggest that the presence of DM in such objects is a plausible hypothesis.
The final Fig. 9(c) illustrates the results of the analysis including all the constraints on mass and radius (I)-(VII). The dominant contribution to the posterior distribution arises from constraints in the region of canonical masses, as well as from the ultracompact object, resulting in an even stronger preference for lower masses of S.
@@
Figure 9: Posterior for an additional set of constraints: (VI) high-mass BW pulsar PSR J0952-0607 (Romani et al., 2022) and (VII) ultracompact HESS J1731-347 (Doroshenko et al., 2022).
@@
Figure 10: Posterior distribution with new NICER constraints (VIII) from the millisecond pulsar PSR J0614–3329 (Mauviard and others, 2025)
@@ 6 Summary and conclusions
Motivated by the theoretical existence of a stable sexaquark configuration, S=uuddss, with a mass below $2054$ MeV, we have explored its role as a bosonic DM candidate within the core of NSs. Using the DD2Y-T model to describe the hadronic phase, incorporating both hyperons and the sexaquark (S) as a bosonic DM component, we modeled the superconducting QM phase with the nlNJL model. A smooth crossover transition between these two phases was constructed using the RIC method. We find that hybrid configurations reach a maximum S DM fraction of about $12$-$15\%$.
Our results show that the inclusion of S particles with a weak repulsive interaction, modeled via a positive mass shift, leads to a softening of the EOS. This effect is consistent with the role of the first established hexaquark, the $d^{*}(2380)$, whose inclusion in the NS EOS has also been shown to soften the EOS and alter the deconfinement threshold. This softening plays a critical role in addressing the tension between stiff hadronic EOSs and tidal deformability constraints, particularly around the canonical $1.4M_{\odot}$ mass range. In contrast to purely hadronic or hybrid models without DM, the resulting hybrid stars, comprising a deconfined QM core and a hypernuclear outer layer enriched with S DM, satisfy all current observational constraints on NS properties.
@@
It is worth noting that without the presence of DM, the considered hadronic (DD2Y-T) and hybrid (RIC_DD2Y-T) models are not consistent with $\Lambda_{1.4}$. In addition, by employing the recently updated NICER M-R results for PSR J0030+0451, we showed that the EOSs including S DM are compatible with the most probable observational values, i.e., the $68\%$ credible region, while the hybrid model without S and the pure hadronic EOS are only consistent with the $95\%$ region.
Furthermore, to place the most stringent constraint on the S mass while accounting for all available observational data, we have incorporated the latest NICER M-R measurements of PSR J0614-3329, which became available prior to the submission of this study. We find that, unlike the EOSs without the S component, all models that include DM favor the $95\%$ credible region, with lighter S particles being particularly preferred. This noteworthy result provides strong support for our model in which S particles are produced inside NSs and thus coexist with other possible internal structures, including hyperons and QM.
@@
To investigate the parameter space of S and constrain its mass more precisely, we performed a BA that incorporates all key observational constraints within a consistent statistical framework. This includes all available M-R measurements by the NICER telescope and the tidal deformability constraint from the GW170817 event by LIGO/Virgo detectors. Additionally, we considered two astrophysical extremes: HESS J1731–347, representing the most compact (lightest) known NS, and PSR J0952–0607, the most massive NS observed to date. These sources provide critical boundary conditions for constraining the EOS across the full density range.
To the best of our knowledge, this study incorporates all currently available and relevant observational data to evaluate the viability of the proposed model. We find that our hybrid model, comprising ordinary nuclear matter, hyperons, S bosonic DM, and deconfined QM, satisfies all observational constraints only when S DM is included. This highlights the essential role of the additional softening and thus the smaller radii around $1.4M_{\odot}$ introduced by the S component. The Bayesian inference yields a posterior distribution that disfavors S particle masses higher than $2000\penalty 10000\ \text{MeV}$ (placing them outside the 95.5% credibility region) and strongly favors $m_{S}\lesssim 1935\penalty 10000\ \text{MeV}$ (within the 68.3% credibility region), suggesting that such a DM candidate may contribute meaningfully to the dense matter composition of NSs.
While our study does not claim that the presence of S particles is required to explain NS properties, it provides compelling evidence that their inclusion leads to the improved consistency of our model with current observational data. Future observations, particularly those sensitive to the thermal and cooling behavior of NSs, may offer further insight into the role of such exotic particles in the interiors of compact stars.
We emphasize that in this work, we have focused on testing the astrophysical viability of the S particle as a DM candidate inside NSs. Our assumption of a small fixed value of $x_{S}$ effectively models a weak coupling to the medium and is consistent with current literature. A DM $S$ state remains viable provided its vacuum couplings to baryons are small, while in-medium effects at sufficiently high densities can still shift its effective mass to influence the EOS. Our present results should be interpreted as the limit of equilibrium configurations. We note that a definitive assessment requires future work that (i) computes weak reaction rates and equilibration timescales involving $S$, (ii) implements a dynamical vector interaction for $S$ beyond a linear mass shift, and (iii) treats possible deconfinement-driven dissociation consistently. These improvements will determine whether the equilibrium compositions assumed here are achieved on astrophysical timescales, but they lie beyond the scope of the present study.
In summary, our analysis shows that the inclusion of DM provides a viable mechanism to soften the stiff hadronic EOSs, such as DD2Y-T, enabling them to remain consistent with modern astrophysical constraints. We stress, however, that this conclusion is model dependent, as softer EOSs may not require, and could even be disfavored by, the presence of DM. In this sense, DM should be regarded as one possible solution among several proposed in the literature to reconcile current tensions between the precise measurements of high mass NSs and the existence of very compact NSs, alongside scenarios such as two families of compact stars (Drago et al., 2016; Drago and Pagliara, 2016; Drago et al., 2014), slowly stable hybrid stars (Mariani et al., 2024; Laskos-Patkos and Moustakidis, 2025), or scenarios involving self-bound hybrid stars and hybrid stars that remain radially stable in the context of slow (or even fast) phase transitions (Zhang et al., 2025).
As a final remark, we note that our analysis is not in contradiction with halo-like DM distributions on galactic scales. Rather, our model describes the distribution of S particles inside NSs under the assumption of non-gravitational BM-DM coupling. The shell-like configuration of S DM at the stellar level can, in principle, coexist with extended DM halo structures on larger astrophysical scales. Thus, our results are not in tension with the DM galactic halo distribution but are complementary to it, providing an additional way to probe this model through multi-messenger observations of NSs.